# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v1

@@ 7.4 Foundation Models assisting for Reinforcement Learning
Reinforcement learning has garnered widespread attention from researchers due to its ability to explore the environment by not requiring extensive annotated data. However, it also faces numerous challenges, such as dealing with long-horizon sequences, reusing experience data, and designing reward functions (Kober et al. (2013)). Foundation models have demonstrated the emergence of common sense reasoning, the ability to propose and sequence sub-goals, and visual understanding. Due to the strong capability of foundation models, many studies aim to leverage the unprecedented capabilities of foundation models to address the challenges faced by reinforcement learning. RobotGPT (Jin et al. (2024)) aims to distill the knowledge of the brain ChatGPT into the mind of a small brain trained with reinforcement learning. At the same time, a plethora of literature investigates the use of foundation models to hierarchically solve challenges like long-horizon problems and reward function design.
Norman (Di Palo et al. (2023)) employs LLMs to decompose tasks into subgoals and utilizes CLIP to identify the completion of each subgoal, serving as a signal generator for sparse rewards. Eureka (Ma et al. (2023b)) utilizes LLM to craft a reward function for five-fingered hand pen spinning. Subsequently, it engages in a cyclic process encompassing reward sampling, GPU-accelerated reward evaluation, and reward reflection to progressively refine its reward outputs. In contrast to Eureka’s self-iteration and sparse reward function design, TEXT2REWARD (Xie et al. (2023a)) incorporates human feedback into the iterative updating of the reward function, yielding a dense reward function.
Following Fig. 7 and Table. 1, pre-trained foundation models can engage with Policy modules in various forms. However, there is currently no benchmark indicating the most effective approach. Regarding end-to-end RFMs, there are still many architectural considerations to make models more interpretable and reliably stable.
@@ 8 Manipulation Data Generation
To propel robots into the era of general manipulation, the acquisition of vast amounts of data is indispensable (Padalkar et al. (2023b)). Gathering real-world data incurs substantial human labor costs. RoboVQA (Sermanet et al. (2023)) adopts a bottom-up data collection approach to gather specific data for fine-tuning VideoCoCa. This bottom-up data collection approach eliminates the need for human scene arrangement and fine-turned VideoCoCa can assist human to decide the next movement in data collection process. While collecting data in simulation presents a more economical solution, it still necessitates human effort in crafting both scene generation and task execution code tailored to specific tasks (Wang et al. (2023a)). The notorious sim-to-real gap issue persists in collecting data in simulation. Therefore, datasets for manipulating deformable and articulated objects still need to be acquired in the real world. Regardless of whether it’s in a real or simulated environment, improving the efficiency of data collection is essential. The mainstream approach is dataset augmentation, primarily divided into object-level and task-level augmentation.
Figure 8: Foundation Models for Manipulation Data Generation. 1) Object-Level Augmentation: The mask of the corresponding object is marked in the image through object segmentation, followed by generative rendering to modify the object’s visual or pose (Wang et al. (2023b)). 2) Task-Level Augmentation: Method 1 generates scene code and policy code by inputting new task instructions into LLM (Wang et al. (2023a)). In simulation scenarios, task execution success can be directly detected through check code, while in real-world scenarios, VLM can be used to detect task success (Ge et al. (2023)). In contrast to Method1, Method2 utilizes LLM to generate new object descriptions, followed by VGM to create new object meshes, forming corresponding scenes. However, the stability of code generation by LLM is comparatively lower. Hence, task instructions are decomposed into subgoals using LLM, followed by RL, Action Primitive, and Gradient Optimization algorithms to generate corresponding actions (Wang et al. (2023b)).
@@
In object-level augmentation, the main idea is to use semantic segmentation to extract masks for each object, and then employ generative rendering methods to alter the object’s visual or pose (Mandlekar et al. (2023); Mandi et al. (2022)). GenAug (Chen et al. (2023c)) leverages language prompts with a generative model to modify object textures and shapes, adding new distractors and background scenes. ROSIE (Yu et al. (2023)) localizes the augmentation region with an open vocabulary segmentation model and then runs image editor to perform text-guided image editing.
In task-level augmentation, the main idea is to use LLMs to generate new task objectives or manually input new task objectives. VGMs is then used to generate meshes for the corresponding task scene objects, or LLMs directly generate simulation code for the corresponding task scene. VLMs or LLMs generate check code to determine the success of executing the new task. GenSim (Wang et al. (2023a)) leverages the code generation capability of LLMs to produce simulation task codes based on new task descriptions. RoboGen (Wang et al. (2023b)) surpasses GenSim in autonomy. It autonomously utilizes LLM to create task descriptions based on specific robot and object inputs and then uses VGM to generate 3D meshes based on these descriptions. Ha et al. (2023) decomposes the new task description into sub-goals using LLM, and then the plan is grounded into a sequence of 6 DOF exploration primitives. An LLM infers a success function code snippet and uses it to verify and label the task as succeeded or failed. PAFF (Ge et al. (2023)) notes that incorrect actions also correspond to new tasks and uses fine-tuned CLIP to label the erroneous trajectories with appropriate task labels. DIAL (Xiao et al. (2022a)) uses fine-tuned CLIP as automatic labelers for robot data, effectively importing internet-scale knowledge into existing datasets with limited ground truth annotations.
Following Fig. 8, most current foundation models for manipulation data generation focus on image rendering or simulation scenes. Exploring the use of foundation models in real-world scenarios to assist in real-time data labeling is also worthwhile.
@@ 9 Discussion
In this survey, we aim to outline the opportunities brought by foundation models for a general-purpose robot framework. We believe the potential of embedding foundation models into manipulation tasks as a viable path towards achieving general manipulation. However, the primary applications of LLMs, VFMs, VLMs, LMMs and VGMs focus only on certain aspects of general operational capabilities, such as reasoning, perception, multimodal understanding, and data generation. The current framework for RFMs resembles human learning, requiring no calibration or ontology parameters, directly learning the mapping between vision and action. However, this demands extensive data for learning, posing a crucial issue of constructing a data close-loop, and ensuring safety remains an unresolved concern. Therefore, this paper proposes a framework of robot learning for manipulation towards achieving general manipulation capabilities and detailing how foundation models can address challenges in each module of the framework. However, there are still many open questions in this survey. In this section, we delve into several open questions that we are particularly concerned about.
@@ 9.1 What is the framework for general manipulation?
The ultimate universal manipulation framework should be capable of interacting with humans to manipulate arbitrary objects in open-world scenarios and achieve diverse operational objectives. However, the interaction between agent and human involves not only recognizing intentions but also learning new skills or improving old skills from human experts in the external world. Open-world scenarios can be static or dynamic, with objects being either rigid or deformable, and task objectives ranging from short-horizon to long-horizon. We designate the restriction of the agent’s learning capability to improving old skills and manipulating rigid objects in static scenes to achieve short-horizon task objectives as Level 0 (L0), the current algorithm implementation has a high probability of achieving L0 product deployment. However, safety and accuracy remain paramount concerns.
@@
From the definition and considerations of safety and accuracy at the L0 level, this paper proposes a framework for a general manipulation capability. Given that the scenarios are static, the framework is designed in a modular, sequential manner. To facilitate module migration, it is preferable for each module to be plug-and-play. Given the current reliance on human-in-the-loop mechanisms in autonomous driving and medical robotics to ensure safety, this framework aims for human-agent interaction through corrective instructions to ensure the safety of manipulation actions. The corrective action can be collected into the dataset and then improve old skills through offline training.
@@
During agent operation, continuous human supervision is not always feasible. Hence, integrating real-time monitoring through parallel surveillance videos during robot execution could enhance safety. The framework in this paper does not explicitly denote this parallel safety monitoring module, as it resembles the post-conditions detection module. The post-conditions detection module analyzes the robot’s operation video to identify reasons for task failure, facilitating post-hoc correction to ensure task success. If the algorithm’s task execution safety is 80%, and the monitoring module predicts safety at 80% as well, the probability of risky movements reduces to 4%. Of course, for household robots, ensuring an over 99% safety rate is imperative. Initially, transitioning nanny from in-home caretakers to cloud-based monitoring of household robot operations appears to be the safest approach and a nanny oversees multiple household robots, allowing human intervention to promptly halt or correct hazardous behaviors. Later, by gathering extensive data to improve model accuracy. This framework still has shortcomings, such as the sequential execution order leads to inefficient characteristics.
@@ 9.3 What kind of learning capability should a general manipulation framework possess?
As an intelligent agent for general manipulation, it is inevitable that one cannot learn all the skills of an open-world in one go, hence possessing a certain learning capability is necessary. Within the framework of this paper, a module of corrective instruction is introduced, enabling the robot to rectify its actions. These corrective demonstrations are incorporated into the manipulation dataset and used to improve the policy offline through fine-tuning. However, this approach still focuses on learning old task skills and cannot acquire new ones.
@@
The model of Policy should possess the capability of interactive few-shot continue online learning to acquire a new skill and reinforce the policy’s mastery of the new skill through corrective instruction offline. Interactive refers to the ability to learn through human demonstration or by observing instructional videos. Learning through human demonstration often requires physical control or teleoperation, which is less natural. Learning through observation of instructional videos aligns better with human learning patterns. However, when humans learn from teachers, they often do not predict the teacher’s trajectory but rather understand the high-level description of the actions, akin to VLaMP (Patel et al. (2023)). Few-shot continue learning enables the agent to learn new skills with minimal demonstrations without forgetting previously learned skills. Online learning entails processing observed data instantly and enabling the model to learn as quickly as possible.
@@
This paper categorizes models trained using manipulation datasets as RFMs. RFMs significantly enhance the policy’s generalization. For instance, LEO (Huang et al. (2023a)) directly outputs target pose, while RoboFlamingo (Li et al. (2023b)) directly outputs delta pose. Directly outputting target pose is easier to train compared to delta pose, although delta pose better aligns with human behavior. Currently, RFMs are fine-tuned using pre-trained models, but this deprives them of self-exploratory learning. Designing an architecture for RFMs is crucial. Utilizing foundation models to aid reinforcement learning still holds many unexplored areas, such as applying VLMs to dense reward function.

### v2

@@ 7.4 Foundation Models assisting for Reinforcement Learning
Reinforcement learning has garnered widespread attention from researchers due to its ability to explore the environment by not requiring extensive annotated data. However, it also faces numerous challenges, such as dealing with long-horizon sequences, effectively exploring, reusing experience data, and designing reward functions (Kober et al. (2013)). Foundation models have demonstrated the emergence of common sense reasoning, the ability to propose and sequence sub-goals, and visual understanding. Due to the strong capability of foundation models, many studies aim to leverage the unprecedented capability of foundation models to address the challenges faced by reinforcement learning. RobotGPT (Jin et al. (2024)) aims to distill the knowledge of the brain ChatGPT into the mind of a small brain trained with reinforcement learning. At the same time, a plethora of literature investigates the use of foundation models to hierarchically solve challenges like long-horizon problems, effectively exploring and reward function design.
Norman (Di Palo et al. (2023)) employs LLMs to decompose tasks into subgoals and utilizes CLIP to identify the completion of each subgoal, serving as a signal generator for sparse rewards. Eureka (Ma et al. (2023b)) utilizes LLM to craft a reward function for five-fingered hand pen spinning. Subsequently, it engages in a cyclic process encompassing reward sampling, GPU-accelerated reward evaluation, and reward reflection to progressively refine its reward outputs. In contrast to Eureka’s self-iteration and sparse reward function design, TEXT2REWARD (Xie et al. (2023a)) incorporates human feedback into the iterative updating of the reward function, yielding a dense reward function. FAC (Ye et al. (2023a)) proposes using knowledge from foundation models as policy prior knowledge to improve sampling efficiency, as value prior knowledge to measure the values of states and as success-reward prior knowledge to provide final feedback on task success.
Following Fig. 7 and Table. 1, pre-trained foundation models can engage with Policy modules in various forms. However, no benchmark indicates which approach is the most effective at present. Regarding end-to-end RFMs, there are still many architectural considerations to make models more interpretable and reliably stable.
@@ 8 Manipulation Data Generation
To propel robots into the era of general manipulation, the acquisition of vast amounts of data is indispensable (Padalkar et al. (2023b)). Collecting real-world data requires a lot of human labor and expensive remote teleoperation equipment. There are currently two methods for data collection: the bottom-up approach and the top-down step-by-step approach. The bottom-up approach focuses on having the robot perform trajectories first. Then, it uses methods like crowd-sourcing to label the data. The top-down step-by-step approach involves decision-makers setting task labels. The robot then performs tasks according to these labels. RoboVQA (Sermanet et al. (2023)) shows that the bottom-up approach is more efficient in data collection compared to the top-down step-by-step approach. It uses the collected data to train VideoCoCa, which helps in handling long-horizon tasks with a hierarchical method. DIAL (Xiao et al. (2022a)) uses a fine-tuned CLIP to replace humans in labeling robot trajectories during bottom-up data collection. This transforms the robot manipulation dataset on the internet into the robot-language manipulation dataset. PAFF (Ge et al. (2023)) points out that incorrect robot trajectories can be linked to new tasks and uses fine-tuned CLIP to label the incorrect robot trajectories with appropriate task labels.
Generating lots of data in simulation is a cheaper solution. However, it still requires human effort to create both scene generation and task execution code for specific tasks (Wang et al. (2023a)). Moreover, the notorious sim-to-real gap issue remains a challenge in transferring policies trained in simulation to real-world applications. But there are many methods to address the sim-to-real challenge. Matas et al. (2018) trains the policy fully in simulation through domain randomization and then successfully deployed in the real world, even though it has never encountered real deformable objects. Therefore, simulation plays an important role in manipulation and this section will analyze existing simulators, scene generation, demonstration generation and sim-to-real gap challenge. Regardless of whether it’s in a real or simulated environment, improving the efficiency of the existing dataset is essential. The mainstream approach is dataset augmentation.
8.1 Simulator
The current mainstream simulators (Zhou et al. (2023)) include PyBullet (Coumans and Bai (2016)), MuJoCo (Todorov et al. (2012)), CoppeliaSim (Rohmer et al. (2013)), NVIDIA Omniverse and Unity. Pybullet is easy to use and integrate, but its graphics are quite basic. It is not suitable for applications that require complex visual effects. Therefore, Pybullet is often used together with Blender (Shi et al. (2024)). Mujoco offers a high-precision physics engine. It is suitable for simulating articulated and deformable object manipulation. However, it has a high entry barrier for beginners. CoppeliaSim offers a wide range of ready-made environments, objects, and prototyping robotic systems for users. However, when dealing with many robots or complex scenes, CoppeliaSim may encounter performance issues. NVIDIA Omniverse provides real-time physics simulation and realistic rendering. However, it requires significant computational resources. NVIDIA Omniverse offers many interfaces. Users can use these to develop various applications. For example, Issac Gym is a platform for robot reinforcement learning, developed using Omniverse. Unity offers rich visual effects and a user-friendly interface. It allows for the creation of highly interactive applications. However, its physics engine is still not precise enough. The basic components of a simulator are the physics engine and the renderer. Improvements in these components can enhance the capability of sensors in simulations, such as optical tactile sensors (Chen et al. (2023d)). We also hope the simulators can add sound engine and other features. This would make the simulated world feel more real. Learning-based simulators also show great potential. For example, Sora (Brooks et al. (2024)) and UniSim (Yang et al. (2023b)) use vast amounts of data from the internet to simulate the visual effects of many different actions.
8.2 Scene and Demonstration Generation
To create simulation scenes for manipulation tasks, two methods can be used. Real-to-Sim method converts real scenes to simulation. Task-specific automation method generates task-specific simulation scenes automatically. Real-to-Sim method can accurately mimic the real world, but it limits the diversity of scenes. The task-specific automation method can create more diverse scenarios. This approach increases the variety of collected demonstrations.
The Real-to-Sim method needs to reconstruct the real scene (in Sec. 6) and to determine the positions of objects. 3D meshes of objects can be made from point cloud data or from RGB images using a generative model (Chen et al. (2024b)). But this Real-to-Sim method for generating object 3D meshes still fails to capture the object’s material and physical properties accurately such as liquid object.
The task-specific automation method requires specific assets tailored to the task. These assets must be imported and sized appropriately, with the scene also being configured to suit the task. RoboGen (Wang et al. (2023b)) utilizes LLM to generate relevant assets, asset sizes, asset configuration, scene configuration based on the task proposals and use text-to-image-to-3D generation to create the corresponding assets. These assets are imported into the simulator to generate the appropriate scene. Finally, using VLM for task-specific scene verification. GenSim (Wang et al. (2023a)) uses LLMs to generate new task and task scenario codes based on the pre-cached scene codes in a task library.
To collect demonstrations in simulations, different approaches can be used based on task complexity. For simple tasks, like a two-finger gripper picking up a cube, a hard-coding method (Wang et al. (2022b)) can be used. However, for more complex tasks, remote teleoperation (Chen et al. (2024a)) or skill library (Ha et al. (2023)) should be employed. Building skill library can be done using reinforcement learning or gradient optimization methods. RoboGen (Wang et al. (2023b)) shows that gradient-based trajectory optimization is better for fine-grained manipulation tasks with soft bodies, like shaping dough into a specific form. On the other hand, reinforcement learning and evolutionary strategies are more effective for contact-rich tasks and continuous interactions with other components in the scene.
8.3 Sim-to-Real Gap Solutions
The sim-to-real problem is a widespread issue across machine learning, not limited to manipulation (Zhao et al. (2020)). The goal is to successfully transfer the policy from the simulation (source domain) to the real world (target domain). The gap in the manipulation tasks between the simulation and the real-world includes two main types: visual gap and dynamic gap. Visual gap refers to the difference between the vision information produced by the renderer and the vision information in the real world. The dynamic gap consists of several factors. First, there is a difference between the physics engine used in simulations and real-world physics. Second, the properties of objects, including robots, contribute to the object dynamic gap. Lastly, there is a control gap in robots, such as variations in static errors caused by different PID parameters. Currently, there are three main approaches to address sim-to-real gap: system identification, domain randomization, and transfer learning (Zhao et al. (2020)).
System identification (Kristinsson and Dumont (1992)) aims to create an accurate mathematical model for a physical system to make the simulator more realistic. However, it is impossible to accurately build models of complex environments in simulators.
Domain randomization (Ramos et al. (2019)) involves adding random disturbances to the parameters in simulation. This can include various elements, generally divided into visual and dynamic randomization. Visual randomization covers visual parameters like lighting, object textures, and camera positions. Dynamic randomization covers dynamic parameters like object sizes, surface friction coefficients, object masses, and actuator force gains. By experiencing diverse simulated environments, the policy can adapt to a broad range of real-world conditions. For the policy, the real world is essentially just another disturbed environment. However, parameter randomization requires human expertise. Ma et al. (2024) demonstrates that LLM excels in selecting randomized parameters and determining the randomization distribution. This makes domain randomization more automated.
Transfer learning (Yu and Wang (2022); Tan et al. (2018)) involves using limited real-world data to adapt a policy trained on a abundant simulation data to the real world. Treat policies in the real-world and in the simulation as different tasks. We can use task transfer methods for transfer learning. For example, Rusu et al. (2017) uses the progressive network to apply knowledge from a policy trained in simulation to a new policy trained with limited real-world data, without losing the previous knowledge. Treat the policies in the real-world and in the simulation as the same task, even though the data distributions differ. We can use domain adaptation methods to address this issue. Three common methods for domain adaptation are discrepancy-based (Lyu et al. (2024)), adversarial-based (Eysenbach et al. (2020)), and reconstruction-based methods (Bousmalis et al. (2016)). Discrepancy-based methods measure the feature distance between the source and target domains using predefined statistical metrics. This helps to align their feature spaces. Adversarial-based methods use a domain classifier to determine whether features come from the source or target domain. Once trained, the extractor can produce features that are invariant across both domains. Reconstruction-based methods also aim to find shared features between domains through setting up an auxiliary reconstruction task and using the shared features to recover the original input.
The methods discussed above assume that the target domain remains unchanged. However, many physical parameters of the same robot can change significantly. Factors like temperature, humidity, positioning, and wear and tear over time can all affect these parameters. This makes it harder to bridge the sim-to-real gap. To address this issue, DORA (Zhang et al. (2024b)) uses an information bottleneck principle. It aims to maximize the mutual information between the dynamics encoding and environmental data. At the same time, it minimizes the mutual information between the dynamics encoding and the behavior policy actions. Transic (Jiang et al. (2024)) proposes a data-driven approach that enables successful sim-to-real transfer using a human-in-the-loop framework.
8.4 Dataset Augmentation
@@
Current dataset augmentation methods primarily operate at the object level. The main idea is to use semantic segmentation to extract masks for each object, and then employ generative rendering methods to alter the object’s visual or pose (Mandlekar et al. (2023); Mandi et al. (2022)). GenAug (Chen et al. (2023c)) leverages language prompts with a generative model to modify object textures and shapes, adding new distractors and background scenes. ROSIE (Yu et al. (2023)) localizes the augmentation region with an open vocabulary segmentation model and then runs image editor to perform text-guided image editing.
Following Fig. 8, LLMs can generate credible descriptions or code for task scenes. VGMs produce 3D object meshes and render textures. Nonetheless, the validity of the generated task scenes must be ultimately assessed by VLMs.
Figure 8: Foundation Models for Manipulation Data Generation. Current mainstream simulators include Pybullet, MuJoCo, CoppeliaSim, NVIDIA Omniverse, and Unity. Meanwhile, learning-based generative models used as simulators have shown potential. Simulation environment generation can be classified into Real-to-Sim and Task-specific automation methods. In the Real-to-Sim method, assuming the object’s position is known, the main challenge lies in constructing the object’s 3D mesh. This can be achieved through scanning technique or by using VGM to generate the 3D mesh directly from RGB image (Chen et al. (2024b)). In the Task-specific automation method, LLM can output scene descriptions or scene code based on task instruction. When the output is a scene description, VGM generates the objects and arranges them according to the description. Meanwhile, the generated scene need to be evaluated by VLM. (Wang et al. (2023b)). When the output is scene code, it directly generates the corresponding scene (Wang et al. (2023a)). However, this requires substantial prior knowledge of scene code within the task library. There are three methods for generating demonstrations in a scene: Hard-code, Teleoperation, and Skill Library. When building skill library, gradient optimization is effective in training skill for deformable tasks and reinforcement learning works better for contact-rich tasks (Wang et al. (2023b)). Solutions for the Sim-to-Real gap include System Identification, Domain Randomization, and Transfer Learning. For data augmentation, VFM is used to segment images first, and then VGM renders the object’s texture on the masked image.
@@ 9 Discussion
In this survey, we aim to outline the opportunities brought by foundation models for general manipulation. We believe the potential of embedding foundation models into manipulation tasks as a viable path towards achieving general manipulation. However, the primary applications of LLMs, VFMs, VLMs, LMMs and VGMs focus only on certain aspects of general manipulation capability, such as reasoning, perception, multimodal understanding, and data generation. The current framework for RFMs resembles human learning, requiring no calibration or robotic parameters, directly learning the mapping between observations and action. However, this demands extensive data for learning, posing a crucial issue of constructing a data close-loop, and ensuring over a 99% success rate remains an unresolved concern. Therefore, this paper proposes a framework of robot learning for manipulation towards achieving general manipulation capability and detailing how foundation models can address challenges in each module of the framework. However, there are still many open questions in this survey. In this section, we delve into several open questions that we are particularly concerned about.
@@ 9.1 What is the framework for general manipulation?
The ultimate general manipulation framework should be able to interact with human or other agent and to manipulate arbitrary objects in open-world scenarios and achieve diverse manipulation tasks. However, the interaction between robot and human involves not only recognizing intentions but also learning new skills or improving old skills from human experts in the external world. Open-world scenarios may be static or dynamic. Objects can be either rigid or deformable. Task objectives can vary from short-term to long-term. Furthermore, tasks may necessitate different degrees of precision with respect to contact points and applied forces/torques. We designate the restriction of the robot’s learning capability to improving old skills and to manipulating rigid objects in static scenes in order to achieve short-horizon task objectives with low precision requirements for contact points and forces/torques as Level 0 (L0), the current research has a high probability of achieving L0. However, safety and accuracy remain paramount concerns.
@@
Based on the definition and considerations of safety and accuracy at the L0 level, this paper proposes a framework for a general manipulation capability. Given that the scenarios are static, the framework is designed in a modular, sequential manner. To facilitate module migration, it is preferable for each module to be plug-and-play. Given the current reliance on human-in-the-loop mechanisms in autonomous driving and medical robotics to ensure safety, this framework aims for human-robot interaction through corrective instruction to ensure the safety of manipulation actions. The corrective action can be collected into the dataset and then improve old skills through offline training.
@@
During robot execution, continuous human supervision is not always feasible. Hence, integrating real-time monitoring through parallel surveillance videos during robot execution could enhance safety. The framework in this paper does not explicitly denote this parallel safety monitoring module, as it resembles the post-conditions detection module. The post-conditions detection module analyzes the robot’s execution video to identify reasons for task failure, facilitating post-hoc correction to ensure task success. If the algorithm’s task execution safety is 80%, and the monitoring module predicts safety at 80% as well, the probability of risky movements reduces to 4%. Of course, for household robots, ensuring an over 99% safety rate is imperative. Initially, cloud-based monitoring of multiple robots by a single operator, with human intervention to correct erroneous behaviors, appears to be the best approach. This strategy not only reduces labor requirements but also ensures safety. Later, by gathering extensive data to improve model accuracy. This framework still has shortcomings, such as the sequential execution order leads to inefficient characteristics.
@@ 9.3 What kind of learning capability should a general manipulation framework possess?
As an intelligent robot for general manipulation, it is inevitable that one cannot learn all the skills of an open-world during offline development, hence possessing a certain learning capability is necessary. Within the framework of this paper, a module of corrective instruction is introduced, enabling the robot to rectify its actions. These corrective demonstrations are incorporated into the manipulation dataset and used to improve the policy offline through fine-tuning. However, this approach still focuses on learning old task skills and cannot acquire new ones.
@@
The model of Policy should possess the capability of interactive, few-shot, continue, online learning to acquire a new skill and reinforce the policy’s mastery of the newly learned skill through corrective instruction offline. Interactive refers to the ability to learn through human demonstration or by observing instructional videos. Learning through demonstration often requires physical control or teleoperation, which is less natural. Learning through observation of instructional videos aligns better with human learning patterns. However, when humans learn from teachers, they often do not predict the teacher’s trajectory but rather understand the high-level description of the actions, akin to VLaMP (Patel et al. (2023)). Few-shot continue learning enables the robot to learn new skills with minimal demonstrations without forgetting previously learned skills. Online learning entails processing observed data instantly and enabling the model to learn as quickly as possible.
@@
This paper categorizes foundation models trained using manipulation datasets as RFMs. RFMs significantly enhance the policy’s generalization. For instance, LEO (Huang et al. (2023a)) directly outputs target pose, while RoboFlamingo (Li et al. (2023c)) directly outputs delta pose. Directly outputting target pose is easier to train compared to delta pose, although delta pose better aligns with human behavior. Currently, RFMs are fine-tuned using pre-trained models, but this deprives them of self-exploratory learning. Designing an architecture for RFMs is crucial. Utilizing foundation models to aid reinforcement learning still holds many unexplored areas, such as using foundation models to assist robots in reinforcement learning within real-world environments.